OpenStudy (anonymous):

Find f. f '(t) = 6 cos (t) + sec^2 (t), −π/2 < t < π/2, f(π/3) = 3

OpenStudy (freckles):

this requires integration

OpenStudy (freckles):

or if you don't know that word antiderivatives!

OpenStudy (freckles):

what is the antiderivative of cos(t) and sec^2(t)?

OpenStudy (anonymous):

sin(t) + C and tan(t) + C

OpenStudy (freckles):

\[f'(t)=6 \cos(t)+\sec^2(t) \\ f(t)=6 \sin(t)+\tan(t)+C\] then we can agree on this right?:

OpenStudy (freckles):

now just apply f(pi/3)=3 to find C

OpenStudy (anonymous):

I'm not sure how to do that

OpenStudy (freckles):

f(pi/3)=3 meants replace f(t) with 3 and replace t with pi/3

OpenStudy (freckles):

\[3=6 \sin(\frac{\pi}{3})+\tan(\frac{\pi}{3})+C\] then solve for the constant C

OpenStudy (anonymous):

I got C = 0.40 (rounded to 2 decimal places)

OpenStudy (freckles):

when you solved for C?

OpenStudy (freckles):

how did you get that

OpenStudy (anonymous):

yes, exact answer is -3Root3 / 2

OpenStudy (anonymous):

[(-3Root3) / 2 ]+3 I meant

OpenStudy (freckles):

something looks a little fishy about that you know sin(pi/3)=sqrt(3)/2 and cos(pi/3)=1/2 so tan(pi/3)=sqrt(3) right?

OpenStudy (freckles):

\[3=6 \sin(\frac{\pi}{3})+\tan(\frac{\pi}{3}) +C \\ 3=6 \frac{\sqrt{3}}{2}+\sqrt{3}+C \\ 3=\frac{6}{2} \sqrt{3}+\sqrt{3}+C \\ 3=3 \sqrt{3}+\sqrt{3}+C\] how many sqrt(3)'s do you see on the right hand side

OpenStudy (anonymous):

2

OpenStudy (freckles):

I see 4.

OpenStudy (anonymous):

ah 3 + 1

OpenStudy (freckles):

\[3 \sqrt{3}+\sqrt{3} \\ \sqrt{3}+\sqrt{3}+\sqrt{3}+\sqrt{3} \\ 4 \sqrt{3}\] yep yep :)

OpenStudy (freckles):

\[3=4 \sqrt{3}+C\]

OpenStudy (freckles):

now try solving for C

OpenStudy (anonymous):

-4sqrt3 + 3

OpenStudy (freckles):

tons cute :)

OpenStudy (freckles):

Now they were looking for your f that satisfied that given condition so just replace out C earlier with this number you found

OpenStudy (freckles):

our*

OpenStudy (freckles):

\[f(t)=6\sin(t)+\tan(t)+C\] that is replace this C with -4 sqrt(3)+3 and you are done

OpenStudy (anonymous):

so f(x) = 6sin(t) + tan(t) + -4sqrt(3) + 3